A person invested $20 000 for 9 years at 7.7% a interest compounded monthly. Calculate the future value of the investment.
A) $45793. 72 B) $38991.62 C) $39905.67 D) $33860.00

Answer by stanbon(75887)

A person invested $20 000 for 9 years at 7.7% a interest compounded monthly. Calculate the future value of the investment.
-------------
A(t) = P(1+(r/n))^(n*t)
A(9) = 20,000(1+0.077/12))^(12*9)
A(9) = 20,000*1.0064^108
A(9) = 20,000*1.9953
A(9) = $39905.67
--------------------
Cheers,
Stan H.
--------------
A) $45793. 72 B) $38991.62 C) $39905.67 D) $33860.00